OpenStudy (anonymous):

a ball of mass 0.5 kg is rolling across a table top with a speed of 5 m/s. when the ball reaches the edge it rolls down a incline plane onto the floor 1.0 meters below. what is the speed of the ball when it reaches the floor?

OpenStudy (xishem):

Use the law of conservation of energy:\[KE_i+U_i=KE_f+U_f\]

OpenStudy (anonymous):

yeah I know that. but how exactly do I do the problem?

OpenStudy (xishem):

By defining the kinetic and potential energies at the beginning and end and solving the equation for the final velocity. Try defining each quantity separately first.

OpenStudy (anonymous):

???

OpenStudy (xishem):

What is the formula for the kinetic energy of an object?

OpenStudy (anonymous):

KE=1/2mv^2

OpenStudy (xishem):

And for the potential energy of an object?

OpenStudy (anonymous):

PE=mgh

OpenStudy (xishem):

Ok, so use the law of conservation of energy by replacing those terms with those equations.

OpenStudy (anonymous):

I'm sure you think I'm stupid but I haven't slept in days and a friend of mine commited suicide a few days ago. so really physics is the last thing on my mind but it has to be done

OpenStudy (xishem):

Still, this isn't a place to bring homework for someone to do for you. If you are willing to accept help and hints, that's okay. But I (and many others) won't do your homework for you.

OpenStudy (anonymous):

didn't ask for you to

OpenStudy (anonymous):

actually I've done half of it myself but am stuck on this one question so lay off

OpenStudy (xishem):

Then provide what you've done so far and what you are stuck on, and I'm happy to help.

OpenStudy (anonymous):

PE= mgh> 0.5 (9.8)(1.0)= PE=4.9

OpenStudy (xishem):

Ok. So that's the initial potential energy of the ball. What about the initial kinetic energy?

OpenStudy (anonymous):

I don't know. I'm confused

OpenStudy (anonymous):

6.25

OpenStudy (anonymous):

I think

OpenStudy (xishem):

Ok. The kinetic energy is this: \[KE=\frac{1}{2}mv^2=\frac{1}{2}(0.5kg)(5\frac{m}{s})^2=6.25J\]Yep! Good job. Now try and find the potential energy after the ball has rolled down the slope.

OpenStudy (anonymous):

wtf is that!?!?

OpenStudy (xishem):

What do you mean?!

OpenStudy (anonymous):

I'm trying to read the big long jumble

OpenStudy (xishem):

It's just finding the kinetic energy of the object. Go ahead and try to find the potential energy after the ball rolls down the slope.

OpenStudy (anonymous):

oh wait.. never mind. I get it

OpenStudy (anonymous):

the whole fraction bracket thing threw me off

OpenStudy (anonymous):

won't it just be the same as the original?

OpenStudy (xishem):

Well, the mass and gravity will be the same, but the height will have changed. If the initial height was at 1 meter, and it falls down 1 meter, then it will be at some new height. You'll have to calculate the potential energy using that height.

OpenStudy (anonymous):

it says the ball rolls down an incline onto the floor 1.0 meter below without bouncing and I am to calculate the speed of the ball when it reaches the floor

OpenStudy (xishem):

Yes. You can calculate the speed by knowing the potential/kinetic energy before and after the fall, using:\[KE_i+U_i=KE_f+U_f\]You've found the initial energies, now you just need to find the final potential energy. Then you will be able to use that formula to find the velocity.

OpenStudy (anonymous):

but how do I know what the other height is??

OpenStudy (xishem):

You used 1 meter to calculate your first potential energy, right? And since the ball dropped 1 meter off of the able, the new height will be 0 meters.

OpenStudy (anonymous):

please don't judge me. I usually catch on quickly but I'm exhausted

OpenStudy (xishem):

That's fine. Here, let me draw a picture. One sec.

OpenStudy (anonymous):

so it's 0?

OpenStudy (xishem):

Yes. When you're using gravitational potential energy, you can set any height to be your 0 height. In this case, by saying that the initial potential energy was from 1 meter, that means that 1 meter below that would be 0 meters.

OpenStudy (anonymous):

but when you multiply the mass+gravity by 0 it just comes out to be 0

OpenStudy (xishem):

That's right. It is 0. Now let's look back at the conservation of energy equation:\[KE_i+U_i=KE_f+U_f\]\[6.25J+4.91J=\frac{1}{2}mv_f^2\]Do you see where to go from here?

OpenStudy (anonymous):

what is this i+u?

OpenStudy (xishem):

Ah. Is the LaTeX not rendering correctly for you? The above should look like this:

OpenStudy (anonymous):

we don't use anything with an i

OpenStudy (xishem):

It just means: initial kinetic energy + initial potential energy = final kinetic energy + final potential energy It's only to be able to differentiate between the two.

OpenStudy (anonymous):

so I basically add 4.9 and 6.25

OpenStudy (xishem):

You have to solve that equation for the velocity. Adding 4.9 and 6.25 will only tell you the kinetic energy when the ball gets to the floor, not the velocity.

OpenStudy (anonymous):

well yeah I knew that. I meant that would be my next step

OpenStudy (xishem):

Yep. That's right. And then you just have to solve that equation to get your final answer.

OpenStudy (anonymous):

11.15=1/2(0.5)v^2 > v=sqrt of 2(9.8)(0.5)

OpenStudy (anonymous):

that's wrong.

OpenStudy (xishem):

\[11.15J=\frac{1}{2}mv^2\]\[\sqrt{\frac{(2)11.15J}{0.5kg}}=6.68\frac{m}{s^2}\]That's what I get.

OpenStudy (anonymous):

I've got 2 more to do and no clue how to do them

OpenStudy (xishem):

I'd be glad to help you, but I have to get off for the night. Good luck!

OpenStudy (anonymous):

well it has to be done tonight